# DMITRY SHIROKOV

### v1

@@ 1 Introduction
The Clifford geometric algebras can be regarded as unified language of mathematics 6, 22, 18, physics 19, 11, 10, 36, engineering 8, and computer science 12, 7. The Clifford geometric algebras are isomorphic to the classical matrix algebras. In particular, the complexified Clifford geometric algebras $\mathcal{G}^{\mathbb{C}}_{p,q}:=\mathbb{C}\otimes\mathcal{G}_{p,q}$ are isomorphic to the following complex matrix algebras:
@@
The paper is organized as follows. In Section 2, we discuss real and complexified geometric algebras (GA) and introduce the necessary notation. In Section 3, we discuss an operation of Hermitian conjugation in GA, introduce a positive scalar product, a norm, unitary space and unitary groups in GA. Also we discuss faithful representations of GA and present an explicit form on one of them. In Section 4, we discuss singular value decomposition of multivectors in GA. In Section 5, we discuss a realization of the determinant and other characteristic polynomial coefficients in GA. In Section 6, we introduce a notion of rank of multivector in GA and prove a number of properties of this notion. We prove that this notion does not depend on the choosing of matrix representation and present another equivalent definition of this notion using only GA operations. Examples for cases of small dimensions are presented. In Section 7, we consider the special case of normal multivectors, for which rank can be determined more simply. The conclusions follow in Section 8.
@@ 2 Real and Complexified Geometric Algebras
Let us consider the real Clifford geometric algebra $\mathcal{G}_{p,q}$ 19, 20, 11, 32 with the identity element $e\equiv 1$ and the generators $e_{a}$, $a=1,2,\ldots,n$, where $n=p+q\geq 1$. The generators satisfy the conditions
$e_{a}e_{b}+e_{b}e_{a}=2\eta_{ab}e,\qquad\eta=(\eta_{ab})={\rm diag}(\underbrace{1,\ldots,1}_{p},\underbrace{-1,\ldots,-1}_{q})$
@@
where we have a sum over arbitrary multi-index $A$ of length from 00 to $n$. The projection of $M$ onto the subspace $\mathcal{G}^{k}_{p,q}$ is denoted by $\langle M\rangle_{k}$.
@@
Let us consider the complexified Clifford geometric algebra $\mathcal{G}_{p,q}^{\mathbb{C}}:=\mathbb{C}\otimes\mathcal{G}_{p,q}$ 32. An arbitrary element of $M\in\mathcal{G}^{\mathbb{C}}_{p,q}$ has the form
@@
In addition to the grade involution and reversion, we use the operation of complex conjugation, which takes complex conjugation only from the coordinates $m_{A}$ and does not change the basis elements $e_{A}$:
@@ 3 Hermitian conjugation and unitary groups in Geometric Algebras
Let us consider an operation of Hermitian conjugation $\dagger$ in $\mathcal{G}^{\mathbb{C}}_{p,q}$ (see 21, 32):
@@
$(M,M)\geq 0,\quad(M,M)=0\Leftrightarrow M=0.$ (10)
Using this scalar product we introduce inner product space over the field of complex numbers (unitary space) in $\mathcal{G}^{\mathbb{C}}_{p,q}$.
@@
$||M||:=\sqrt{(M,M)}=\sqrt{\langle M^{\dagger}M\rangle_{0}}.$ (11)
@@
Let us present an explicit form of one of these representations of $\mathcal{G}^{\mathbb{C}}_{p,q}$ (we use it also for $\mathcal{G}_{p,q}$ in 25 and for $\mathcal{G}^{\mathbb{C}}_{p,q}$ in 33). We denote this fixed representation by $\beta^{\prime}$. Let us consider the case $p=n$, $q=0$. To obtain the matrix representation for another signature with $q\neq 0$, we should multiply matrices $\beta^{\prime}(e_{a})$, $a=p+1,\ldots,n$ by imaginary unit $i$. For the identity element, we always use the identity matrix $\beta^{\prime}(e)=I_{N}$ of the corresponding dimension $N$. We always take $\beta^{\prime}(e_{a_{1}a_{2}\ldots a_{k}})=\beta^{\prime}(e_{a_{1}})\beta^{\prime}(e_{a_{2}})\cdots\beta^{\prime}(e_{a_{k}})$. In the case $n=1$, we take $\beta^{\prime}(e_{1})={\rm diag}(1,-1)$. Suppose we know $\beta^{\prime}_{a}:=\beta^{\prime}(e_{a})$, $a=1,\ldots,n$ for some fixed odd $n=2k+1$. Then for $n=2k+2$, we take the same $\beta^{\prime}(e_{a})$, $a=1,\ldots,2k+1$, and
@@
where $\dagger$ is the Hermitian transpose of a matrix. Using the linearity, we get that Hermitian conjugation of matrix is consistent with Hermitian conjugation of corresponding multivector:
@@
Note that the same is not true for an arbitrary matrix representations $\beta$ of the form (12). It is true the matrix representations $\gamma=T^{-1}\beta^{\prime}T$ obtained from $\beta^{\prime}$ using the matrix $T$ such that $T^{\dagger}T=I$.
@@
which we call a unitary group in $\mathcal{G}^{\mathbb{C}}_{p,q}$. Note that all the basis elements $e_{A}$ of $\mathcal{G}_{p,q}$ belong to this group by the definition.
@@ 4 Singular Value Decomposition in Geometric Algebras
31 For an arbitrary multivector $M\in\mathcal{G}^{\mathbb{C}}_{p,q}$, there exist multivectors $U,V\in{\rm U}\mathcal{G}^{\mathbb{C}}_{p,q}$, where
@@ 5 Determinant and other characteristic polynomial coefficients in Geometric Algebras
Let us consider the concept of determinant 24, 5 and characteristic polynomial 25 in geometric algebra. Explicit formulas for characteristic polynomial coefficients are discussed in 1, 2, applications to Sylvester equation are discussed in 26, 27, the relation with noncommutative Vieta theorem is discussed in 28, 29, applications to calculation of elementary functions in geometric algebras are discussed in 3.
@@
The presented algorithm and formulas (20), (21), (26) are actively used to calculate inverse in GA 4, 17, 23.
@@ 6 Rank in Geometric Algebras
${\rm rank}(M)={\rm rank}(\widehat{M})={\rm rank}(\widetilde{M})={\rm rank}(\overline{M})$ (32)
$\qquad={\rm rank}(M^{\dagger})={\rm rank}(M^{\dagger}M)={\rm rank}(MM^{\dagger}),\qquad\forall M\in\mathcal{G}^{\mathbb{C}}_{p,q}.$ (33)
@@
Let us prove ${\rm rank}(M)={\rm rank}(\widetilde{M})$. We have the following relation between the reversion (or the superposition of reversion and grade involution) and the transpose (see 35, 33):
@@
${\rm rank}(A)=N\quad\Leftrightarrow\quad C_{(N)}\neq 0;$ (35)
${\rm rank}(A)=k\in\{1,\ldots,N-1\}\,\Leftrightarrow\,C_{(k)}\neq 0,\,\,C_{(j)}=0,\,j=k+1,\ldots,N;$ (36)
${\rm rank}(A)=0\quad\Leftrightarrow\quad A=0.$ (37)
@@
$C_{(N)}(M^{\dagger}M)$ $=$ $-{\rm Det}(M^{\dagger}M)=-{\rm Det}(M^{\dagger}){\rm Det}(M)$
$=$ $-({\rm Det}M)^{2}=(C_{(N)}(M))^{2},$
$C_{(1)}(M^{\dagger}M)$ $=$ $N\langle M^{\dagger}M\rangle_{0}=N||M||^{2},$
@@
${\rm rank}(M)=\begin{cases}N,\quad&\mbox{if $C_{(N)}(M)\neq 0$,}\\ N-1,\quad&\mbox{if $C_{(N)}(M)=0$ and $C_{(N-1)}(T)\neq 0$,}\\ N-2\qquad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=0$ and}\\ &\mbox{$C_{(N-2)}(T)\neq 0$,}\\ \cdots&\\ 2,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=\cdots=C_{(3)}(T)=0$ and}\\ &\mbox{$C_{(2)}(T)\neq 0$,}\\ 1,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=\cdots=C_{(2)}(T)=0$ and}\\ &\mbox{$M\neq 0$,}\\ 0,\quad&\mbox{if $M=0$.}\end{cases}$ (44)
@@
Let us consider the $\bigtriangleup$-operation 25
@@ 7 The Case of Normal Multivectors
Hermitian multivectors $M^{\dagger}=M$, anti-Hermitian multivectors $M^{\dagger}=-M$, unitary multivectors $M^{\dagger}M=e$ are the particular cases of normal multivectors. For example, the basis elements $e_{A}$ are unitary by the definition. Note that all unitary multivectors have rank equal to $N$.
@@
${\rm rank}(M)=\begin{cases}N,\quad&\mbox{if $C_{(N)}(M)\neq 0$,}\\ N-1,\quad&\mbox{if $C_{(N)}(M)=0$ and $C_{(N-1)}(M)\neq 0$,}\\ N-2\qquad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=0$ and}\\ &\mbox{$C_{(N-2)}(M)\neq 0$,}\\ \cdots&\\ 2,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=\cdots=C_{(3)}(M)=0$ and}\\ &\mbox{$C_{(2)}(M)\neq 0$,}\\ 1,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=\cdots=C_{(2)}(M)=0$ and}\\ &\mbox{$M\neq 0$,}\\ 0,\quad&\mbox{if $M=0$.}\end{cases}$ (50)
@@
We use that normal matrix is always diagonalizable.
@@ 8 Conclusions
In this paper, we implement the notion of rank of multivector in complexified Clifford geometric algebras without using the corresponding matrix representations. Theorem 6.11 involves only operations in geometric algebras. To obtain the results, we use the fact that rank of the matrix $A\in{\rm Mat}(N,\mathbb{C})$ is the number of nonzero (positive) singular values of $A$, that is, the number of nonzero (positive) eigenvalues of the matrix $A^{\dagger}A$. Theorems 6.3, 6.5, 6.11, 7.1 are new. New explicit formulas (47), (49) for the cases of dimensions $3$ and $4$ can be used in various applications of geometric algebras in physics, engineering, and computer science.

### v2

@@ 1 Introduction
The Clifford geometric algebras can be regarded as unified language of mathematics 6, 24, 18, physics 19, 11, 10, 38, engineering 8, and computer science 12, 7, 20. The Clifford geometric algebras are isomorphic to the classical matrix algebras. In particular, the complexified Clifford geometric algebras $\mathcal{G}^{\mathbb{C}}_{p,q}:=\mathbb{C}\otimes\mathcal{G}_{p,q}$ are isomorphic to the following complex matrix algebras:
@@
The paper is organized as follows. In Section 2, we discuss real and complexified geometric algebras (GA) and introduce the necessary notation. In Section 3, we discuss an operation of Hermitian conjugation in GA, introduce a positive scalar product, a norm, unitary spaces, and unitary groups in GA. Also we discuss faithful representations of GA and present an explicit form on one of them. In Section 4, we discuss singular value decomposition of multivectors in GA. In Section 5, we discuss a realization of the determinant and other characteristic polynomial coefficients in GA. In Section 6, we introduce the notion of rank of multivector in GA and prove a number of properties of this notion. We prove that this notion does not depend on the choosing of matrix representation and present another equivalent definition of this notion using only GA operations. Examples for cases of small dimensions are presented. In Section 7, we consider the special case of normal multivectors, for which the rank can be determined more simply. The conclusions follow in Section 8.
@@ 2 Real and Complexified Geometric Algebras
Let us consider the real Clifford geometric algebra $\mathcal{G}_{p,q}$ 19, 22, 11, 34 with the identity element $e\equiv 1$ and the generators $e_{a}$, $a=1,2,\ldots,n$, where $n=p+q\geq 1$. The generators satisfy the conditions
$e_{a}e_{b}+e_{b}e_{a}=2\eta_{ab}e,\qquad\eta=(\eta_{ab})={\rm diag}(\underbrace{1,\ldots,1}_{p},\underbrace{-1,\ldots,-1}_{q}).$
@@
where we have a sum over all multi-indices of length from 00 to $n$. The projection of $M$ onto the subspace $\mathcal{G}^{k}_{p,q}$ is denoted by $\langle M\rangle_{k}$.
@@
Let us consider the complexified Clifford geometric algebra $\mathcal{G}_{p,q}^{\mathbb{C}}:=\mathbb{C}\otimes\mathcal{G}_{p,q}$ 34. An arbitrary element of $M\in\mathcal{G}^{\mathbb{C}}_{p,q}$ has the form
@@
In addition to the grade involution and reversion, we use the operation of complex conjugation, which takes complex conjugation only from the coordinates $m_{A}$ and does not change the basis elements $e_{A}$ of $\mathcal{G}_{p,q}$:
@@ 3 Hermitian conjugation and unitary groups in Geometric Algebras
Let us consider an operation of Hermitian conjugation $\dagger$ in $\mathcal{G}^{\mathbb{C}}_{p,q}$ (see 23, 34):
@@
$(M,M)\geq 0,\quad(M,M)=0\Longleftrightarrow M=0.$ (10)
Using this scalar product we introduce an inner product space over the field of complex numbers (a unitary space) in $\mathcal{G}^{\mathbb{C}}_{p,q}$.
@@
$\|M\|:=\sqrt{(M,M)}=\sqrt{\langle M^{\dagger}M\rangle_{0}}.$ (11)
@@
Let us present an explicit form of one of these representations of $\mathcal{G}^{\mathbb{C}}_{p,q}$ (we use it also for $\mathcal{G}_{p,q}$ in 27 and for $\mathcal{G}^{\mathbb{C}}_{p,q}$ in 35). We denote this fixed representation by $\beta^{\prime}$. Let us consider the case $p=n$, $q=0$. To obtain the matrix representation for another signature with $q\neq 0$, we should multiply matrices $\beta^{\prime}(e_{a})$, $a=p+1,\ldots,n$ by imaginary unit $i$. For the identity element, we always use the identity matrix $\beta^{\prime}(e)=I_{N}$ of the corresponding dimension $N$. We always take $\beta^{\prime}(e_{a_{1}a_{2}\ldots a_{k}})=\beta^{\prime}(e_{a_{1}})\beta^{\prime}(e_{a_{2}})\cdots\beta^{\prime}(e_{a_{k}})$. In the case $n=1$, we take $\beta^{\prime}(e_{1})={\rm diag}(1,-1)$. Suppose we know $\beta^{\prime}_{a}:=\beta^{\prime}(e_{a})$, $a=1,\ldots,n$ for some fixed odd $n=2k+1$. Then for $n=2k+2$, we take the same $\beta^{\prime}(e_{a})$, $a=1,\ldots,2k+1$, and
@@
where $\dagger$ is the Hermitian transpose of a matrix. Using the linearity, we get that Hermitian conjugation of a matrix is consistent with Hermitian conjugation of the corresponding multivector:
@@
Note that the same is not true for an arbitrary matrix representations $\beta$ of the form (12). It is only true for matrix representations $\gamma=T^{-1}\beta^{\prime}T$ obtained from $\beta^{\prime}$ using matrices $T$ such that $T^{\dagger}T=I$.
@@
which we call a unitary group in $\mathcal{G}^{\mathbb{C}}_{p,q}$. Note that all the basis elements $e_{A}$ of $\mathcal{G}_{p,q}$ belong to this group by definition.
@@ 4 Singular Value Decomposition in Geometric Algebras
32, 33 For an arbitrary multivector $M\in\mathcal{G}^{\mathbb{C}}_{p,q}$, there exist multivectors $U,V\in{\rm U}\mathcal{G}^{\mathbb{C}}_{p,q}$, where
@@ 5 Determinant and other characteristic polynomial coefficients in Geometric Algebras
Let us consider the concept of determinant 26, 5 and characteristic polynomial 27 in geometric algebra. Explicit formulas for characteristic polynomial coefficients are discussed in 1, 2, applications to Sylvester equation are discussed in 28, 29, the relation with the noncommutative Vieta theorem is discussed in 30, 31, applications to calculation of elementary functions in geometric algebras are discussed in 3.
@@
The presented algorithm and formulas (20), (21), (26) are actively used to calculate inverse in GA 4, 17, 25. See also 21.
@@ 6 Rank in Geometric Algebras
${\rm rank}(M)={\rm rank}(\widehat{M})={\rm rank}(\widetilde{M})={\rm rank}(\overline{M})={\rm rank}(M^{\dagger})={\rm rank}(M^{\dagger}M)={\rm rank}(MM^{\dagger}),\qquad\forall M\in\mathcal{G}^{\mathbb{C}}_{p,q}.$ (32)
@@
Let us prove ${\rm rank}(M)={\rm rank}(\widetilde{M})$. We have the following relation between the reversion (or the superposition of reversion and grade involution) and the transpose (see 37, 35):
@@
${\rm rank}(A)=N\quad$$\Longleftrightarrow$ $\quad C_{(N)}\neq 0;$ (34)
${\rm rank}(A)=k\in\{1,\ldots,N-1\}\quad$$\Longleftrightarrow$ $\quad C_{(k)}\neq 0,\quad C_{(j)}=0,\quad j=k+1,\ldots,N;$ (35)
${\rm rank}(A)=0\quad$$\Longleftrightarrow$ $\quad A=0.$ (36)
@@
$C_{(N)}(M^{\dagger}M)$ $=$ $-{\rm Det}(M^{\dagger}M)=-{\rm Det}(M^{\dagger}){\rm Det}(M)=-({\rm Det}M)^{2}=(C_{(N)}(M))^{2},$
$C_{(1)}(M^{\dagger}M)$ $=$ $N\langle M^{\dagger}M\rangle_{0}=N\|M\|^{2},$
@@
${\rm rank}(M)=\begin{cases}N,\quad&\mbox{if $C_{(N)}(M)\neq 0$,}\\ N-1,\quad&\mbox{if $C_{(N)}(M)=0$ and $C_{(N-1)}(T)\neq 0$,}\\ N-2\qquad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=0$ and $C_{(N-2)}(T)\neq 0$,}\\ \cdots&\\ 2,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=\cdots=C_{(3)}(T)=0$ and $C_{(2)}(T)\neq 0$,}\\ 1,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(T)=\cdots=C_{(2)}(T)=0$ and $M\neq 0$,}\\ 0,\quad&\mbox{if $M=0$.}\end{cases}$ (43)
@@
Let us consider the $\bigtriangleup$-operation 27
@@ 7 The Case of Normal Multivectors
Hermitian multivectors $M^{\dagger}=M$, anti-Hermitian multivectors $M^{\dagger}=-M$, unitary multivectors $M^{\dagger}M=e$ are the particular cases of normal multivectors. For example, the basis elements $e_{A}$ of $\mathcal{G}_{p,q}$ are unitary by the definition. Note that all unitary multivectors have rank equal to $N$.
@@
${\rm rank}(M)=\begin{cases}N,\quad&\mbox{if $C_{(N)}(M)\neq 0$,}\\ N-1,\quad&\mbox{if $C_{(N)}(M)=0$ and $C_{(N-1)}(M)\neq 0$,}\\ N-2\qquad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=0$ and $C_{(N-2)}(M)\neq 0$,}\\ \cdots&\\ 2,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=\cdots=C_{(3)}(M)=0$ and $C_{(2)}(M)\neq 0$,}\\ 1,\quad&\mbox{if $C_{(N)}(M)=C_{(N-1)}(M)=\cdots=C_{(2)}(M)=0$ and $M\neq 0$,}\\ 0,\quad&\mbox{if $M=0$.}\end{cases}$ (49)
@@
We use that a normal matrix is always diagonalizable.
@@ 8 Conclusions
In this paper, we establish the notion of rank of multivector in complexified Clifford geometric algebras without using the corresponding matrix representations. Theorem 6.11 involves only operations in geometric algebras. To obtain the results, we use the fact that rank of the matrix $A\in{\rm Mat}(N,\mathbb{C})$ is the number of nonzero (positive) singular values of $A$, that is, the number of nonzero (positive) eigenvalues of the matrix $A^{\dagger}A$. Theorems 6.3, 6.5, 6.11, 7.1 are new. New explicit formulas (46), (48) for the cases of dimensions $3$ and $4$ can be used in various applications of geometric algebras in physics, engineering, and computer science.
@@
[Acknowledgements] The author is grateful to the anonymous reviewers for their careful reading of the paper and helpful comments on how to improve the presentation.
{ack}